OpenStudy (anonymous):

MEDAL + FAN Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (acxbox22):

you can do any simple one like x^2+x+8 which i think cant be factored but can be by the quad. formula

OpenStudy (anonymous):

2t^2-5t+11t^2 t(2t^2+11t-5)=0 then do quadratic formula -11+- (square root sign here) 121-4(2)(5)/2(2) -11+-(sq root here) 81/4

OpenStudy (anonymous):

a= 2t^2 b=11t c=5

OpenStudy (anonymous):

I know it can only be solved by quadratic formula because when you reach step: t(2t^2+11t-5)=0 and try to factor it doesnt work because no numbers multiply to 10 and add to 11

OpenStudy (acxbox22):

that equation wont work because t(2t^2+11t-5)=0 simplifies to 2t^3+11t^2-5t which is a cubic polynomial